OpenStudy (anonymous):

initial value problem y' = xy +x^3 y(0) = -1

OpenStudy (anonymous):

y(0)=-1 y'=0(-1)+0^3 y'=0

OpenStudy (anonymous):

y' - xy=x^3 IF = e^(-xdx) multiply by IF on both sides

OpenStudy (anonymous):

d(ye^(-x^2 / 2)) = x^3 * e^(-x^2 / 2)dx

OpenStudy (anonymous):

now someone remind me how to integrate the rhs..im growing old

OpenStudy (amistre64):

gummage, that is not appropriate language for this forum; please keep it clean

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

take x^2 = u and then do integral by parts

OpenStudy (anonymous):

you mean u = x^3?

OpenStudy (anonymous):

was this integration

OpenStudy (anonymous):

u=x^2/2 du= xdx RHS becomes ue^(-u)du

OpenStudy (anonymous):

as it is x^3(e^(-x^2/2))

OpenStudy (anonymous):

it becomes ue^(-u)du/2 sorry

OpenStudy (anonymous):

now use integral by parts

OpenStudy (mathteacher1729):

After integrating and using the initial condition, you obtain: \[y(x) = -x^2+e^{x^2/2}-2\] Attached is a picture of the slope field with the initial condition drawn in for fun. :) http://www.math.rutgers.edu/~sontag/JODE/JOdeApplet.html

OpenStudy (anonymous):

he needs to learn to do himself....u guys are extremely dependent on websites...

OpenStudy (anonymous):

did u do it gummage?

OpenStudy (anonymous):

how does the x^3 change into u/2?

OpenStudy (mathteacher1729):

NOTES: http://tutorial.math.lamar.edu/Classes/DE/IntroFirstOrder.aspx http://ocw.mit.edu/courses/mathematics/18-03-differential-equations-spring-2010/readings/ VIDEOS: http://www.khanacademy.org/#differential-equations http://ocw.mit.edu/courses/mathematics/18-03-differential-equations-spring-2010/video-lectures/ In particular "First order ODEs" seem to be right up your ally. :) If you have windows, download WINPLOT, it is awesome and free. http://math.exeter.edu/rparris/winplot.html A great way to visualize your work. HIM1628-- I agree, sometimes people just "Give answers". I try whenever possible to give helpful resources and guidance. It seemed that this discussion broke down the problem fairly well, so here are some final thoughts on ODEs. :)

OpenStudy (anonymous):

see gummage havent u got\[\int\limits_{}^{}x^{3}e^{-x^{2}/2}dx\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now \[x^{2}/2=u\] then xdx=du

OpenStudy (anonymous):

yep got that

OpenStudy (anonymous):

now u can write the RHS as \[\int\limits_{}^{}2ue^{-u}du\]

OpenStudy (anonymous):

xdx becomes du x^2 becomes 2u the divide by 2 ws my bad

OpenStudy (anonymous):

ive got x^4*e^-u du

OpenStudy (anonymous):

sorry got it now. thanks